Can someone help me on this problem?
I need to solve this literal equation for the indicated variable.
2x + 3y = 6 )for x) indicated variable is linear equation
Thanks,
Sher

Answer by checkley71(8403)

2X+3Y=6 OR 2X=6-3Y OR X=6/2-3Y/2 OR X=3-3Y/2